OpenStudy (anonymous):

The area of a circle is found using the formula A=nr^2 , where r is the radius. if the area of a circle is 36n square feet, what is the radius, in feet? N= Pi

OpenStudy (jdoe0001):

$$ Area = \pi r^2\\ 36\pi = \pi r^2\\ \text{solve for "r"}\\ \cfrac{36\pi}{\pi} = r^2 $$

OpenStudy (jim766):

A = pi *r^2 just fill in what they gave you 36pi = pi r^2 just need to solve for r, -- --- pi pi 36 = r^2

OpenStudy (jim766):

what is the sq rt of 36?

OpenStudy (jdoe0001):

so, to isolate "r", get \(\large \sqrt{}\) on both sides, what do you get?

OpenStudy (anonymous):

sq rt of 36 is 6?

OpenStudy (jdoe0001):

so, r = 6

OpenStudy (jim766):

right that is the radis

OpenStudy (anonymous):

6n or just 6?

OpenStudy (jim766):

errrr 6 ft

OpenStudy (anonymous):

in the answer choice there is 6n (pi) and just 6 ? which is it..

OpenStudy (jdoe0001):

6 units, whatever units, meters, feet, who knows, \(\pi\) for this case, is just a CONSTANT used, \(\pi = 3.1416\)

OpenStudy (anonymous):

soooo which one do i use

OpenStudy (jim766):

it is just 6 ft

OpenStudy (anonymous):

okat thanks both of yall so much :)) can you help me with my next question?

OpenStudy (jdoe0001):

from $$ Area = \pi r^2\\ 36\pi = \pi r^2\\ \text{solve for "r"}\\ \cfrac{36\pi}{\pi} = r^2 $$

OpenStudy (jim766):

36(3.14) = 3.14(6^2) check it 113.04 = 113.04

OpenStudy (jdoe0001):

notice, the \(\pi\) above and below, will cancel out, so, as Jim766 said, you're only left with "36" all by its lonesome

OpenStudy (anonymous):

yes